OpenStudy (aaronandyson):

A wire is used t form a circle with diameter 35mm.The wire is also used in making 5 diameters which divide the circle into 10 equal sectors Find 1. the total length of wire used 2.the area of each sector

OpenStudy (aaronandyson):

@ganeshie8 @Sadworld

OpenStudy (aaronandyson):

@sarahjane116

OpenStudy (aaronandyson):

@butterflydreamer

OpenStudy (mathmath333):

correction total length of wire \((C_1+5(2r_2))\)

OpenStudy (zpupster):

35mm - circumference of circle what is the diameter of the 35 mm circle? hint : C=Pi(diameter) or C = 2(pi)r 35/pi=diam (5*diam) + 35 = length of wire

OpenStudy (aaronandyson):

17.5 times 5 plus 35?

OpenStudy (aaronandyson):

@zpupster 35mm is the diameter given -.-

OpenStudy (aaronandyson):

im really confused :(

OpenStudy (aaronandyson):

@mathmath333

OpenStudy (mathmath333):

let the length of wire be divide in two parts \(x_1~and ~ x_2\) the first part is used to form form a circle with diameter 35mm \((d_1=35)\) so it is equal to its circumference so \(x=\pi d_1=35\pi\) first see if it makes sense

OpenStudy (mathmath333):

i mean \(x_1=35\pi\)

OpenStudy (aaronandyson):

area of the circle is 110?

OpenStudy (mathmath333):

how ?

OpenStudy (aaronandyson):

sorry i meant circumference = 110

OpenStudy (mathmath333):

yes

OpenStudy (aaronandyson):

now?

OpenStudy (mathmath333):

also the wire is used in making 5 diameters which divide the circle into 10 equal sectors let area of each sector be \(A\) so \(\large A_s=\dfrac{area ~of ~circle}{10}\\ =\dfrac{\pi (d_1/2)^2}{10}\\\)

OpenStudy (mathmath333):

also \(x_2=5\times d_1\) so \(total ~wire=x_1+5\times d_1\)

OpenStudy (mathmath333):

see if the answer matches with book

OpenStudy (aaronandyson):

area of each sector?

OpenStudy (mathmath333):

@zpupster 35mm is the length of diameter given

OpenStudy (aaronandyson):

AREA OF EACH SECTOR? AREA OF CIRCLE? IM SO CONFUSED!!!!!!

OpenStudy (zpupster):

sorry

OpenStudy (mathmath333):

its ok @zpupster lol, calm down @AaronAndyson and please read my reference

OpenStudy (aaronandyson):

yet me confused

OpenStudy (mathmath333):

which part ?

OpenStudy (aaronandyson):

is it 22/7*17.5*17.5?

OpenStudy (aaronandyson):

for the full circle

OpenStudy (mathmath333):

yes

OpenStudy (aaronandyson):

9625?

OpenStudy (mathmath333):

its comming \(962.1\) for me

OpenStudy (mathmath333):

area of full circle ^

OpenStudy (mathmath333):

so \(96.21\) area of each sector

OpenStudy (aaronandyson):

back of the book says 96.25

OpenStudy (mathmath333):

lol its \(\approx\) approximate

OpenStudy (mathmath333):

that difference comes because the exact value of \(\pi\) is not \(22/7\) its \(3.14...............\)

OpenStudy (aaronandyson):

?????????????????????????????????????????

OpenStudy (mathmath333):

forgot that lol , any doubt still ?

OpenStudy (aaronandyson):

more questions

OpenStudy (mathmath333):

sry have to go , but when i come i will have a look at it , good luck